OpenStudy (anonymous):

A mobile phone passes through the hands of three persons each of whom loses 5%. If the third person sells it for Rs 6859, What did the first person pay for it? A) Rs 7600 B) Rs 9600 C) Rs 10,000 D) Rs 8000

OpenStudy (nubeer):

its A the next person pays 95% of the previous price. 2nd person= 7600*0.95 3rd person= 7600*0.95*0.95 = 6859